number of atoms in 560 g of fe(atomic mass 56 g mol -1 ) is: options are. 1.Twice that in 70g N 2.half that in 20 g H 3.both 1 and 2 4.none of these

  1. number of atoms in 560 g of fe(atomic mass 56 g mol-1) is:
options are.
1.Twice that in 70g N
2.half that in 20 g H
3.both 1 and 2
4.none of these

Grade:12th pass

2 Answers

Aksh
6 years ago
No of atoms=No of moles×Avagadro noNo of moles=Mass of sample/atomic massNo of moles of Fe = 560/56=10No of Fe atoms=10×No. (No=Avagadro no)No of moles of N =70/14=5No of N atoms=5No not equal to FeNo of moles of H atoms= 20/1=20No of atoms of H =20No not equal to FeTherefore the ans is none of these(d)
Anjali kumari
6 years ago
As mole=given mass/ molecular mass.Therefore, 560g/56g mole`= 56 mole.Now, number of atoms=10×6×10^23= 6×10^24.In option (i) mole=70/14 mole.Therefore number of atoms=2×70×6×10^23/14=6×10^23×10=6×10^24.Similarly In option (ii) number of atoms=6×10^24.Therefore the answer is (iii)